Razz Corporation has budgeted for the following sales:

July $425,000
August $510,000
September     $605,000

October

$860,000
November $715,000
December $680,000

Sales are collected as follows: 10% in the month of sale; 60% in the month following the sale; and the remaining 30% in the second month following the sale.

1. What is the budgeted total cash collected in December?

2. In Razz's budgeted balance sheet at December 31, at what amount will accounts receivable be shown?
Answer #1
Concepts and reason

Financial Statement: The basic financial statement includes the balance sheet which shows the entities overall assets, liabilities and the net worth as on date, income statement which shows the net income generated by the entity over a specified time period and the cash flow statements which shows the cash outflows and inflows over a period of time from all the entries.

Balance sheet: It is the financial statement which includes the assets, equity and liability of a corporation at a particular time. It identifies the company’s financial position. The steps for preparations of Balance Sheet are given below:

1. The first heading in the balance sheet preparation is assets. Under which the balances of currents assets are presented first, then the balances of long-term investments are entered and then the balances of plant, property, and equipment are entered.

2. The second heading in the balance sheet preparation is liabilities and shareholder’s equity. The liabilities are presented first, under which the balances of current liabilities and then the long-term debts balances are presented. Stockholder’s equity balances are entered after the liabilities.

3. Ensure that the total assets balance is equal to the sum of total liabilities and shareholder’s equity.

Fundamentals

Assets: It can be defined as the resources which are controlled and owned by the organization and which are capable of providing some future benefits for operating the core business of the organization. Examples of assets are Furniture, Machinery, Car, Cash and many others.

Current Assets: It means the short- term assets which can be convertible into cash within a period of one accounting year or operating cycle. It includes Accounts Receivables, Cash and Cash Equivalents, Inventories, and many others.

Accounts Receivable: It is an asset for the company. It is a consideration, which the company is having the right to receive for the goods or services provided to the customers. It is the payment which the company will receive in the future from its customers.

Cash: It is an asset in liquid form. Whenever any payment is made by the corporation, it can be made in the form of cash and whenever any income or revenue is received by the corporation, cash balance increases.

Calculate the budgeted total cash collected in the month of December.

CashcollectionDecember=10%×SaleDecember+60%×SaleNovember+30%×SaleOctober=(10%×$680,000)+(60%×$715,000)+(30%×$860,000)=$68,000+$429,000+$258,000=$755,000\begin{array}{c}\\{\rm{Cash}}\,{\rm{collectio}}{{\rm{n}}_{{\rm{December}}}} = 10\% \times {\rm{Sal}}{{\rm{e}}_{{\rm{December}}}} + 60\% \times {\rm{Sal}}{{\rm{e}}_{{\rm{November}}}} + 30\% \times {\rm{Sal}}{{\rm{e}}_{{\rm{October}}}}\\\\ = \left( {10\% \times \$ 680,000} \right) + \left( {60\% \times \$ 715,000} \right) + \left( {30\% \times \$ 860,000} \right)\\\\ = \$ 68,000 + \$ 429,000 + \$ 258,000\\\\ = \$ 755,000\\\end{array}

Calculate the receivables at the end of December month to be presented on the company’s balance sheet.

AccountsReceivableDecember=SaleNovember+SaleDecember=(30%×$715,000)+(90%×$680,000)=$214,500+$612,000=$826,500\begin{array}{c}\\{\rm{Accounts}}\,{\rm{Receivabl}}{{\rm{e}}_{{\rm{December}}}} = \,{\rm{Sal}}{{\rm{e}}_{{\rm{November}}}} + {\rm{Sal}}{{\rm{e}}_{{\rm{December}}}}\\\\ = \left( {30\% \times \$ 715,000} \right) + \left( {90\% \times \$ 680,000} \right)\\\\ = \$ 214,500 + \$ 612,000\\\\ = \$ 826,500\\\end{array}

Ans:

The cash collected in the December month is $755,000.

Accounts Receivable at the end of December is $826,500.